# Lightweight Deep Learning-Based Channel Estimation for RIS-Aided Extremely Large-Scale MIMO Systems on Resource-Limited Edge Devices

### v1

@@ VII Simulation
Figure 4: Performance comparison between a $16\times 32$ patch size and the results obtained with a $32\times 32$ patch size after doubling the RIS dimensions to $32$.
Fig. 4 compares the NMSE performance of the proposed method using different patch shapes with that of the B-LMMSE, LMMSE, and LS methods. Each point in the figure represents the average NMSE over 2,000 test samples. The results demonstrate that the proposed architecture significantly outperforms the LS method, particularly at low SNR values, where it exhibits substantially better accuracy. Moreover, B-LMMSE and LMMSE initially perform similarly to the proposed method, but the performance gap becomes more significant as the SNR increases toward 0 dB. Beyond this point, the LMMSE method begins to follow the performance trend of LS, while the B-LMMSE performs worse than LS due to the binary reflection strategy, where only one element is active at a time, resulting in insufficient received power for accurate channel estimation. Additionally, a slight variation in NMSE is observed between different patch sizes. Specifically, the $16\times 32$ patch consistently achieves higher NMSE compared to the $32\times 32$ patch. This improvement stems from the increased RIS dimensions, which enable slightly better extraction of spatial correlations between the RIS elements and BS antennas. The NMSE variation between the above patch shapes ranges from approximately 2% to 15%. Notably, the difference is more pronounced at low SNR levels, where the impact of noise is higher, and it gradually decreases as the SNR increases.
Figure 5: Compared performance by breaking each data point into patches of size $32\times 32$ for testing and evaluating it against using the entire data point without patching.
Fig. 5 presents a comparative analysis of training performance with and without patches during the testing phase. The motivation for employing patches in testing is to evaluate the prediction performance on patches of the same size as those used in training $32\times 32$ versus prediction on the data point of size $1024\times 128$. This allows us to assess the impact of patch-based training when applied to patch-based versus full-resolution testing. As observed in Fig. 5, there is a slight improvement in NMSE when testing with patches compared to testing without patches. The improvement ranges from 0.1% to 8%, with the highest improvement occurring at low SNR values. Specifically, a peak improvement of 8% is observed at $-5$ dB SNR. This result can be attributed to training on patches, enabling the model to estimate noise more effectively at a finer level in low SNR conditions. However, while testing with patches provides a slight performance improvement over testing without patches, it comes at the cost of significantly increased computational time. Specifically, testing with patches requires approximately four times more computation, which may pose challenges in low-latency scenarios. This additional time overhead stems from dividing data into patches, performing individual predictions, and recombining them to estimate the channel.
@@
Fig. 6 presents a performance comparison of the direct channel estimation from the user to BS. Similar to the cascaded channel training, we consider 10,000 samples of the direct link channel, denoted as $\mathbf{b}_{k}\in\mathcal{C}^{M\times 1}$. To make the data suitable for training, we transform this 1D data into a 2D representation of size $32\times 32$, leveraging that these dimensions correspond to the antennas’ horizontal and vertical arrangement. As observed in Fig. 6, there is a significant difference in NMSE between the proposed method and the LS algorithm, particularly at low SNR values. This discrepancy arises due to the high noise levels in low SNR scenarios, which pose a major challenge for LS. However, the proposed algorithm demonstrates superior performance, providing reliable channel estimates even under severe noise conditions. Furthermore, the performance gap between the proposed method and LS decreases as SNR increases. This trend can be attributed to the lower noise levels at higher SNR values, which reduce the estimation error for both methods.
@@
Fig. 7 compares the NMSE performance of correlated and uncorrelated Rayleigh fading channels for different training patch sizes. The correlated channel exhibits the lowest NMSE compared to the uncorrelated Rayleigh fading channel. This is because correlated fading patterns are more structured, allowing deep learning models to capture better and exploit these patterns. However, in practical scenarios, achieving uncorrelated Rayleigh fading is nearly impossible, as there are always dominant directions that carry most of the signal energy. Furthermore, at low SNR levels, the NMSE performance of both correlated and uncorrelated channels is similar due to the overwhelming presence of noise. A noticeable performance difference between correlated and uncorrelated Rayleigh fading channels begins to emerge at an SNR of 0 dB, with an initial gap of approximately 0.5 dB. As the SNR increases, this gap widens significantly, with the correlated channel demonstrating up to a 2.5 dB improvement at an SNR of 15 dB.
@@
Fig. 8 illustrates the impact of increasing the number of phase shift matrices (pilot overhead) on NMSE performance, comparing it with LS and LMMSE methods. As observed, increasing the phase shift matrix positively influences NMSE by reducing estimation error. However, introducing excessive phase shifts leads to significant pilot overhead, which is undesirable in practical communication systems. A sharp decrease in NMSE is noticeable when transitioning from $L=96$ to $L=128$. This improvement is because, as $L$ approaches 128, the system moves closer to achieving a full-rank LS matrix. A full-rank LS matrix ensures better-conditioned channel estimation, leading to effective model training. Beyond $L=128$, NMSE continues to decrease, but at a slower rate, as the benefits of increasing $L$ diminish. Moreover, our proposed method yields good results, while the difference between the LS and LMMSE methods remains minimal, as both rely on the full-rank assumption.
Figure 9: Comparative evaluation of varying the number of RIS $N$ and antennas $M$ on system performance at an SNR of 10 dB.
Fig. 9 illustrates the impact of varying the number of BS antennas $M$ and RIS elements $N$ on NMSE performance compared to LS estimation. As shown, increasing $M$ has a negligible effect on NMSE, even when training with different patch sizes. This is because the training process relies on patches, allowing the model to learn latent representation in the data irrespective of the input size. Consequently, during inference, varying $M$ does not significantly impact the results. A similar trend is observed for N, where the number of RIS elements $N$ is increased. This indicates that training the model with patches enables it to capture underlying data correlations independent of $M$ and $N$. In contrast, traditional schemes are inherently affected by variations in these parameters, highlighting a key advantage of the proposed learning-based approach.
@@ VIII Conclusions
In this work, we addressed the critical challenge of cascaded channel estimation in XL-MIMO systems empowered by RIS. Recognizing the limitations of traditional and existing deep learning approaches in resource-constrained environments, we proposed a novel, lightweight framework tailored for next-generation wireless networks. Using spatial correlation and patch-based training, our solution effectively reduces the dimensionality of the training data while preserving key features. We further enhance performance using a customized UNet 3+ architecture with integrated denoising capabilities, allowing accurate estimation even with limited training data. Extensive evaluations confirm that our framework enhances estimation accuracy and significantly reduces computational overhead, maintaining efficiency even as the number of antennas and RIS elements scales in XL-MIMO systems, making it well-suited for real-time deployment on edge devices. This pioneering approach lays the foundation for scalable, energy-efficient, and practical RIS-enabled XL-MIMO systems in 6G and beyond.

### v2

@@ VI Computational Complexity
VII Simulation
Furthermore, our comprehensive literature survey revealed that, to the best of our knowledge, no prior work has investigated deep learning-based cascaded channel estimation for RIS-assisted XL-MIMO systems. Among existing studies, we identified CDRN [DL1] and SSL [DL4] as the most relevant works. However, these schemes are limited by their reliance on a relatively small number of antennas and RIS elements. Specifically, their training procedures involve processing full-dimensional matrices $\tilde{\mathbf{G}}^{(i)}_{k}\in\mathbb{C}^{M\times N}$ for each data point. Given the scale and complexity of XL-MIMO systems for edge devices, where the number of antennas and RIS elements can reach hundreds or even thousands, training such models on large-scale datasets is computationally and memory intensive, making them impractical for edge deployment. To address these limitations and ensure a fair evaluation, we trained the benchmark models using the same set of patches extracted from the dataset as those used in our model. This strategy aligns with edge device constraints while maintaining a consistent comparison with the proposed method. Moreover, several classical estimation techniques, including LS [LS], LMMSE [LMMSE], and B-LMMSE [BLMMSE], are included as relevant benchmarks. The performance of the proposed framework is evaluated in terms of the normalized mean squared error (NMSE), defined as $\text{NMSE}=\frac{\mathbb{E}\!\left[\|\mathbf{G}-\hat{\mathbf{G}}_{\text{proposed}}\|_{F}^{2}\right]}{\mathbb{E}\!\left[\|\mathbf{G}\|_{F}^{2}\right]}$ where $\hat{\mathbf{G}}_{\text{proposed}}$ and $\mathbf{G}$ denote the estimated and ground-truth channels, respectively.
In this paper, we generated a synthetic dataset based on the system models described in Section III. The dataset was created with consideration for the memory constraints of edge devices. Specifically, a total of 10,000 data samples were generated, and from each sample, a representative data patch was extracted to encapsulate the underlying characteristics of the corresponding data point. As the antenna array has dimensions of $32\times 32$ and the RIS elements are arranged as $16\times 8$, the patch sizes were selected as $16\times 32$ and $32\times 32$ to align with the spatial correlation properties between the antenna and RIS elements. The dataset was split into training and validation sets, with 70% allocated for training and 30% for validation. Additionally, we created a separate testing dataset for different SNRs, with a sample size of 2,000 for each SNR value. Each point in the figures represent the average NMSE over 2,000 test samples.
Figure 5: Performance comparison of the proposed scheme with benchmark methods and with varying patch sizes.
Figure 5 presents the performance comparison between the proposed scheme and benchmark algorithms across an SNR range of -15 dB to 20 dB. At low SNR levels (15-15 dB to 10-10 dB), where noise dominates, the proposed scheme outperforms existing methods, achieving an error reduction of approximately $13.65$ dB compared to LS, $0.8$ dB over LMMSE and BLMMSE, and $0.5$ dB over CDRN and SSL. As the SNR increases to 5-5 dB and 0 dB, the proposed scheme consistently outperforms LS by up to $11.58$ dB, LMMSE and BLMMSE by 6.486.48-$8.29$ dB, and learning-based methods CDRN and SSL by $3.02$ and $2.58$ dB, respectively, demonstrating superior generalization and feature extraction capabilities. In the moderate-to-high SNR region (5–20 dB), the performance advantage of the proposed model persists even as competing methods approach saturation. At 10 dB, it outperforms LS, LMMSE, and BLMMSE by $4.69$, $4.26$, and $13.89$ dB, respectively, and achieves gains of $1.7$ and $1.0$ dB over CDRN and SSL. At 15 dB and 20 dB, the proposed scheme continues to show improvements, achieving NMSE reductions of $2.74$ and $1.18$ dB over LS, 2.692.691.161.16 dB over LMMSE, and 0.560.560.230.23 dB and 0.450.450.100.10 dB over CDRN and SSL, respectively. These results demonstrate that the proposed model maintains its robustness and accuracy even under high-SNR conditions. Additionally, we compare the proposed method’s performance for patch sizes of $16\times 32$ and $32\times 32$. The results indicate that the NMSE varies across patch sizes, with the $32\times 32$ configuration consistently achieving lower NMSE values than the $16\times 32$ case, showing improvements ranging from $0.4$ dB to $0.08$ dB. This improvement arises from the larger patch size capturing stronger spatial correlations, thereby reducing estimation errors (especially at low SNRs), albeit at the cost of higher complexity.
Figure 6: Compared performance by varying the dataset size used for model training.
Figure 6 compares the performance of the proposed framework with benchmark algorithms across training dataset sizes from 3,000 to 12,000 samples. Since LS and LMMSE are analytical estimators that operate independently of training data, their performance remains unchanged, exhibiting NMSE values approximately $6.19$ dB and $5.88$ dB higher than those of the proposed framework, respectively, across all dataset sizes. In contrast, the data-driven models, CDRN and SSL, demonstrate a clear dependence on dataset size. For smaller datasets (3,000–5,000 samples), CDRN shows a performance degradation of about 2.52.52.62.6 dB relative to the proposed scheme, while SSL lags by approximately 0.80.81.181.18 dB. As the dataset size increases (6,000–9,000 samples), both models show gradual improvements due to enhanced feature learning and greater data diversity. Specifically, CDRN’s NMSE decreases from $2.49$ dB to $2.32$ dB, and SSL’s NMSE decreases from $1.41$ dB to $1.32$ dB. However, their performance remains inferior to the proposed framework by more than $2$ dB and $1$ dB, respectively. For larger datasets (10,000–12,000 samples), CDRN and SSL achieve NMSEs of approximately $2.03$ dB and $1.58$ dB, respectively, yet these values remain about $2$ dB and $1.5$ dB higher than those of the proposed model. The proposed framework consistently achieves lower NMSE with increasing dataset size, effectively leveraging larger datasets, capturing complex spatial correlations, and maintaining superior generalization compared with CDRN and SSL. Overall, increasing the dataset size improves estimation accuracy but also raises computational complexity and memory requirements, which must be carefully considered when deploying the model on resource-constrained edge devices.
@@ VII Simulation
Figure 7 illustrates the performance comparison for direct channel estimation from the user to the BS across an SNR range of 15-15 dB to $20$ dB. The proposed framework consistently outperforms all benchmark algorithms across the entire SNR range. The LS estimator exhibits the poorest performance, with NMSE values approximately $16.35$ dB, $12.67$ dB, and $9.41$ dB higher than the proposed framework at 15-15 dB, 10-10 dB, and 5-5 dB, respectively. Even at moderate-to-high SNR levels (02020 dB), LS remains significantly inferior, with the performance gap decreasing from $6.65$ dB at 0 dB to $0.41$ dB at $20$ dB. In contrast, the learning-based models CDRN and SSL achieve closer results but still lag behind the proposed method. Specifically, at extremely low SNRs (15-15 dB to 10-10 dB), CDRN performs $0.73$ dB and $0.8$ dB worse, while SSL trails by $0.41$ dB and $0.46$ dB, respectively. At $10$ dB, CDRN and SSL remain $0.3$ dB and $0.17$ dB below the proposed framework, while at $20$ dB, they nearly converge, with small gaps of $0.11$ dB and $0.1$ dB, respectively. These results indicate that, while CDRN and SSL perform competitively at moderate-to-high SNRs, the proposed framework consistently achieves superior accuracy across all SNR levels, particularly under low-SNR and noise-dominant conditions, demonstrating strong generalization and noise resilience.
Figure 8 presents a comprehensive performance comparison of the proposed framework against benchmark schemes under both correlated and uncorrelated fading conditions for NLoS channels. The proposed models for both correlated and uncorrelated channels consistently outperform the corresponding CDRN and SSL benchmark schemes across all SNR levels. At low SNRs (15-15 dB to 10-10 dB), the proposed correlated model achieves significantly lower error levels, outperforming CDRN and SSL under the same correlation by approximately 0.500.502.562.56 dB and 0.490.492.672.67 dB, respectively, while the uncorrelated configuration attains slightly higher NMSE improvements of 0.700.702.772.77 dB and 0.460.462.792.79 dB. As the SNR increases to moderate levels (5-5 dB to $5$ dB), both correlated and uncorrelated versions of the proposed framework continue to demonstrate superior performance, with the correlated model achieving gains of approximately 2.152.152.212.21 dB over CDRN and 3.553.551.581.58 dB over SSL. These consistent improvements highlight the proposed model’s strong ability to capture complex spatial features and learn robust representations as signal reliability improves. At higher SNRs ($10$ dB to $20$ dB), where most models begin to saturate, the proposed correlated configuration maintains performance advantages of 1.691.690.230.23 dB over CDRN and 1.071.070.100.10 dB over SSL, while the uncorrelated framework achieves corresponding gains of approximately 1.151.150.280.28 dB and 1.201.200.200.20 dB. Overall, the correlated variants consistently outperform across all SNR regimes, demonstrating that spatial correlation enhances the model’s ability to exploit inter-element dependencies, leading to more accurate, stable, and noise-resilient channel estimation.
@@
Figure 9 illustrates the impact of increasing the number of phase shift matrices ($L$) on the NMSE performance, comparing the proposed scheme with LS, LMMSE, BLMMSE, CDRN, and SSL methods. As observed, increasing $L$ significantly improves estimation accuracy up to $L=128$, after which the performance gain becomes marginal. At $L=96$, the proposed method achieves an NMSE improvement of approximately $1.3$ dB over CDRN and $1.7$ dB over SSL, while maintaining a substantial margin of more than $4$ dB compared to the LS and LMMSE estimators. The most notable performance gain occurs at $L=128$, where the proposed scheme surpasses CDRN and SSL by around $2.2$ dB and $1.58$ dB, respectively. This improvement is attributed to better channel conditioning as $L$ approaches a full-rank LS matrix, enhancing the model’s learning capability. Beyond $L=128$, further increases in $L$ yield diminishing returns, as seen from the slower NMSE reduction. For higher values of $L$ ($144$, $160$, and $176$), the NMSE continues to decrease slightly, with the proposed method maintaining gains of approximately 1.21.21.51.5 dB over CDRN and 1.21.21.41.4 dB over SSL, indicating diminishing returns but consistently outperforming data-driven and analytical benchmarks.
@@ VIII Conclusions
In this work, we addressed the critical challenge of cascaded channel estimation in XL-MIMO systems empowered by RIS. Recognizing the limitations of traditional and existing deep learning approaches in resource-constrained environments, we proposed a novel, lightweight framework tailored for next-generation wireless networks. Using spatial correlation and patch-based training, our solution effectively reduces the dimensionality of the training data while preserving key features. We further enhance performance using our proposed architecture with integrated denoising capabilities, allowing accurate estimation even with limited training data. Extensive evaluations confirm that our framework enhances estimation accuracy and significantly reduces computational overhead, maintaining efficiency even as the number of antennas and RIS elements scales in XL-MIMO systems, making it well-suited for real-time deployment on edge devices. This pioneering approach lays the foundation for scalable, energy-efficient, and practical RIS-enabled XL-MIMO systems in 6G and beyond.